# The SRG/eROSITA all-sky survey

### v1

@@ Acknowledgements.
TR and AV acknowledge support from the German Federal Ministry of Economics and Technology (BMWi) provided through the German Space Agency (DLR) under project 50 OR 2112. AV acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 450861021. We thank F. Maccagni and P. Serra for providing MeerKAT continuum contours of Fornax A and MeerKAT HI emission contours of ram pressure stripped galaxies. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The eROSITA data shown here were processed using the eSASS software system developed by the German eROSITA consortium. This research has made use of the NASA/IPAC Extragalactic Database (NED) which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. Partly based on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. ChatGPT has been used to speed up writing some of the Python scripts, e.g., for plotting data. This work made use of astrometry.net.References
@@ Appendix A Comparison of eRASS:2 and eRASS3+eRASS4+eRASS5 images
Essentially, the same features appear for both images here as those discussed in the main text for the full dataset (eRASS:5). In particular, Finger 1–3 and Excess 1–4 are present in both. This demonstrates visually that these surface brightness features are not caused by statistical fluctuations. This is expected given the noise-suppression properties of the wavelet-filtering process.
@@ Appendix B Hard band eRASS:5 image
In Fig. 1 we showed that the Fornax cluster lies in a region with particularly stable background emission; background is low and does not show obvious strong variations. Still, given the scale of several degrees covered by the eROSITA image of the Fornax cluster, it is difficult to exclude emission variations in our Milky Way Galaxy as the underlying cause of the observed surface brightness fluctuations. Furthermore, varying absorption and uncertainties in our correction procedure may also induce apparent surface brightness variations. Both effects, emission from the low temperature plasma in our Galaxy and absorption by its neutral gas, are in practice only important well below a photon energy of 1 keV. Therefore, if one or both of these effects indeed gave rise to the features discussed in this paper, we would expect them to vanish in an image created using only photons with an energy of at least 1 keV. Therefore, we constructed a hard band image using the energy range 1.0–2.3 keV and show it in Fig. 20.
Essentially, the same features appear in this image as in the broad band image. All Fingers and Excess regions are present. Also the Bridge to the Fornax A group becomes particularly clear in this energy band. This demonstrates that these surface brightness features are unlikely to be caused by varying emission or absorption effects in our Galaxy, as these are expected to be very weak above 1.0 keV.
Appendix C Close up views of background clusters
eROSITA can detect extended emission from hot gas not only nearby but out to redshifts even beyond one. In Fig. 15 we showed the distribution of background clusters on top of the eROSITA image of the Fornax cluster and the comparison was discussed in detail in Sect. 3.5. Here, we provide versions of the same image but zoomed into various areas to provide more detail.

### v2

@@ Acknowledgements.
We thank the anonymous referee for carefully reading the manuscript and providing very helpful feedback. TR and AV acknowledge support from the German Federal Ministry of Economics and Technology (BMWi) provided through the German Space Agency (DLR) under project 50 OR 2112. AV acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 450861021. VG acknowledges the financial contribution from the contracts Prin-MUR 2022 supported by Next Generation EU (M4.C2.1.1, n.20227RNLY3 The concordance cosmological model: stress-tests with galaxy clusters. MH acknowledges financial support from the Excellence Cluster ORIGINS which is funded by the DFG under Germany’s Excellence Strategy – EXC 2094 – 390783311. AL acknowledges the support from the National Natural Science Foundation of China (Grant No. 12588202). AL is supported by the China Manned Space Program with grant no. CMS-CSST-2025-A04. We thank F. Maccagni and P. Serra for providing MeerKAT continuum contours of Fornax A and MeerKAT HI emission contours of ram pressure stripped galaxies. This work is based on data from eROSITA, the soft X-ray instrument aboard SRG, a joint Russian-German science mission supported by the Russian Space Agency (Roskosmos), in the interests of the Russian Academy of Sciences represented by its Space Research Institute (IKI), and the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The SRG spacecraft was built by Lavochkin Association (NPOL) and its subcontractors, and is operated by NPOL with support from the Max Planck Institute for Extraterrestrial Physics (MPE). The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg & ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität Munich also participated in the science preparation for eROSITA. The eROSITA data shown here were processed using the eSASS software system developed by the German eROSITA consortium. This research has made use of the NASA/IPAC Extragalactic Database (NED) which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. Partly based on observations obtained with XMM-Newton, an ESA science mission with instruments and contributions directly funded by ESA Member States and NASA. ChatGPT has been used to speed up writing some of the Python scripts, e.g., for plotting data. This work made use of astrometry.net.References
@@
- Coutinho et al. (2022) Coutinho, D., Ramos-Ceja, M. E., Dennerl, K., et al. 2022, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 12181, Space Telescopes and Instrumentation 2022: Ultraviolet to Gamma Ray, ed. J.-W. A. den Herder, S. Nikzad, & K. Nakazawa, 121811A
@@
- Freyberg et al. (2022) Freyberg, M. J., Müller, T., Perinati, E., et al. 2022, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 12181, Space Telescopes and Instrumentation 2022: Ultraviolet to Gamma Ray, ed. J.-W. A. den Herder, S. Nikzad, & K. Nakazawa, 1218155
- Galárraga-Espinosa et al. (2021) Galárraga-Espinosa, D., Aghanim, N., Langer, M., & Tanimura, H. 2021, A&A, 649, A117
- Galárraga-Espinosa et al. (2024) Galárraga-Espinosa, D., Cadiou, C., Gouin, C., et al. 2024, A&A, 684, A63
@@
- Martis et al. (2024) Martis, N. S., Sarrouh, G. T. E., Willott, C. J., et al. 2024, ApJ, 975, 76
- Martizzi et al. (2019) Martizzi, D., Vogelsberger, M., Artale, M. C., et al. 2019, MNRAS, 486, 3766
@@
- Walker et al. (2019) Walker, S., Simionescu, A., Nagai, D., et al. 2019, Space Sci. Rev., 215, 7
- Wang et al. (2024) Wang, W., Wang, P., Guo, H., et al. 2024, MNRAS, 532, 4604
@@ Appendix A Comparison of eRASS:2 and eRASS3+eRASS4+eRASS5 images
Essentially, the same features appear for both images here as those discussed in the main text for the full dataset (eRASS:5). In particular, Fingers 1–3 and Excesses 1–4 are present in both. This demonstrates visually that these surface brightness features are not caused by statistical fluctuations. This is expected given the noise-suppression properties of the wavelet-filtering process.
@@ Appendix B Hard band eRASS:5 image
In Fig. 1 we show that the Fornax cluster lies in a region with particularly stable background emission; background is low and does not show obvious strong variations. Still, given the scale of several degrees covered by the eROSITA image of the Fornax cluster, it is difficult to exclude emission variations in our Milky Way Galaxy as the underlying cause of the observed surface brightness fluctuations. Furthermore, varying absorption and uncertainties in our correction procedure may also induce apparent surface brightness variations. Both effects, emission from the low temperature plasma in our Galaxy and absorption by its neutral gas, are in practice only important well below a photon energy of 1 keV. Therefore, if one or both of these effects indeed gave rise to the features discussed in this paper, we would expect them to vanish in an image created using only photons with an energy of at least 1 keV. Therefore, we constructed a hard band image using the energy range 1.0–2.3 keV and show it in Fig. 18.
Essentially, the same features appear in this image as in the broad band image. All fingers and excess regions are present. Also, the bridge to the Fornax A group becomes particularly clear in this energy band. This demonstrates that these surface brightness features are unlikely to be caused by varying emission or absorption effects in our Galaxy, as these are expected to be very weak above 1.0 keV.
Appendix C Surface brightness profiles in sectors
As discussed at the end of Sect. 3.1, we searched for a large-scale bow shock in various directions. We show the surface brightness profiles of several sectors in Figs. 19 and 20.
Appendix D Close-up views of background clusters
eROSITA can detect extended emission from hot gas not only nearby but out to redshifts even beyond one. In Fig. 13 we show the distribution of background clusters on top of the eROSITA image of the Fornax cluster and the comparison was discussed in detail in Sect. 3.5. Here, we provide versions of the same image but zoomed into various areas to provide more detail.